### v1

# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

Abstract
Recent advancements in Large Language Models (LLMs) have facilitated the development of Multimodal LLMs (MLLMs). Despite their impressive capabilities, MLLMs often suffer from an over-reliance on unimodal biases (e.g., language bias and vision bias), leading to incorrect answers in complex multimodal tasks. To investigate this issue, we propose a causal framework to interpret the biases in Visual Question Answering (VQA) problems. Within our framework, we devise a causal graph to elucidate the predictions of MLLMs on VQA problems, and assess the causal effect of biases through an in-depth causal analysis. Motivated by the causal graph, we introduce a novel MORE dataset, consisting of 12,000 VQA instances. This dataset is designed to challenge MLLMs’ abilities, necessitating multi-hop reasoning and the surmounting of unimodal biases. Furthermore, we propose two strategies to mitigate unimodal biases and enhance MLLMs’ reasoning capabilities, including a Decompose-Verify-Answer (DeVA) framework for limited-access MLLMs and the refinement of open-source MLLMs through fine-tuning. Extensive quantitative and qualitative experiments offer valuable insights for future research. Our project page is at https://opencausalab.github.io/MORE.
Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective
Meiqi Chen [^1] This work was done during her internship at Shanghai AI Laboratory. Peking University meiqichen@stu.pku.edu.cn                        Yixin Cao Singapore Management University caoyixin2011@gmail.com
Yan Zhang [^2] Corresponding author. Peking University zhyzhy001@pku.edu.cn                        Chaochao Lu [^2] Corresponding author. Shanghai AI Laboratory luchaochao@pjlab.org.cn
1 Introduction
Following the success of Large Language Models (LLMs) Ouyang et al. (2022); Touvron et al. (2023b), Multimodal LLMs (MLLMs) OpenAI (2023); Team et al. (2023) have been proposed for various vision-language tasks Fu et al. (2023); Liu et al. (2023b). Despite their promising results, it remains unclear if they truly understand images and text in the context of multi-modal reasoning.
Figure 1: Examples of over-reliance on unimodal biases. MLLMs (e.g., LLaVA) erroneously generate answers due to language bias (indicated by the underlined text below the left image) and vision bias (the right image).
As shown in the knowledge-based Visual Question Answering (VQA) problems of Figure 1, when prompted with “Which country is hosting the next World Cup after this venue?” MLLMs such as GPT-4V OpenAI (2023) and LLaVA Liu et al. (2023a) may capture the language bias of “the next World Cup” and think that the next World Cup will be “the 2022 FIFA World Cup held in Qatar” (which is also outdated knowledge), while ignoring the exact venue presented in the image. Similarly, when presented with an image of “The Shard” in London, MLLM directly identifies “The representative building is The Shard” influenced by vision bias, overlooking the specific constraint “in Berlin” mentioned in the question. These inherent issues pose significant challenges to the reasoning capabilities of MLLMs, particularly when faced with more complex questions.
In order to investigate the issue of MLLMs’ Over-REliance (MORE) on such unimodal biases, we propose a causal framework to interpret and quantify language and vision biases. Specifically, we first define the causal graph of MLLM’s prediction on VQA problems. The causal graph is constructed based on a variety of causal factors that are integral to the prediction process, such as image and question text. Then, we identify a set of interventions in the context of VQA problems, thereby ascertaining the causal effect of unimodal biases on the prediction capabilities of MLLMs via $do$-calculus Pearl (1995). By quantifying such causal effects, we can evaluate MLLM’s sensitivity and robustness against unimodal biases.
Datasets
Knowledge-
based
Multi-hop
Reasoning
Answer Type
Unimodal Biases
Evaluation
Rationale # Size
Visual7W  Zhu et al. (2016) Open-ended 327.9K
VQA (v2) Goyal et al. (2017) Open-ended 1.1M
FVQA Wang et al. (2017) Open-ended 5.8K
OKVQA Marino et al. (2019) Open-ended 14K
S3VQA Jain et al. (2021) Open-ended 7.5K
A-OKVQA Schwenk et al. (2022) Multi-choice 23.7K
INFOSEEK Chen et al. (2023) Open-ended 1.4M
MORE (Ours) Multi-choice 12K
Table 1: Comparison of MORE with other VQA datasets, highlighting its incorporation of external knowledge, multi-hop reasoning, unimodal bias evaluation, and rationale for interpretability.
Based on the above causal analysis, we curate a novel dataset termed MORE, comprising 12,000 VQA instances. This dataset advances existing VQA datasets by introducing a dedicated evaluation of unimodal biases. To facilitate the evaluation, we adopt a Multiple Choice Question (MCQ) format, where each instance consists of an image, a question, and four candidate options. The image is sourced from an existing VQA dataset Chen et al. (2023). For question and option curation, we incorporate a knowledge graph (KG) Wang et al. (2021), allowing us to better simulate MLLMs to navigate the corresponding spurious paths within the causal graph. Specifically, the options consist of one correct answer, and three distractors targeting language bias, vision bias, and multi-hop reasoning, respectively. We also provide the reasoning path, designated as causal rationale, in the KG for each instance, offering interpretability for evaluation. As summarized in Table 1, compared to existing VQA datasets, MORE features external knowledge, multi-hop reasoning, unimodal bias evaluation, and rationale, showing better comprehensiveness. Experimental results on six leading MLLMs reveal that: 1) most MLLMs perform much poorly on MORE, showing an obvious tendency to rely on unimodal biases. 2) MLLMs still struggle to achieve precise semantic understanding when dealing with multimodal reasoning.
Furthermore, to mitigate the unimodal biases and enhance MLLM’s reasoning abilities, we also propose two strategies. The basic idea is to leverage step-by-step multimodal reasoning to mitigate spurious paths. The first strategy involves prompt engineering, featuring limited-access MLLMs (e.g., GPT-4V OpenAI (2023)) through a Decompose-Verify-Answer (DeVA) framework to explicitly guide MLLMs to extract information from multiple modals comprehensively. The second strategy requires fine-tuning open-source MLLMs such as LLaVA Liu et al. (2023a) based on our dataset, instructing them with the generated causal rationale.
Overall, our main contributions are as follows:
  • We propose a causal framework with a well-defined causal graph to interpret and quantify the biases in VQA problems.
  • We construct a new MORE dataset, designed to challenge MLLMs to overcome reliance on unimodal biases via multi-hop causal reasoning, demonstrating superior comprehensiveness compared to existing VQA datasets.
  • We propose two strategies to mitigate the unimodal biases and enhance the reasoning ability of current MLLMs, including prompting and fine-tuning ways.
2 A Causal Framework
Figure 2: Causal graph of MLLM’s Prediction on VQA problems. We highlight the difference between a cognitively-inspired correct reasoning path ($G_{h}$) and the undesired effects that some factors might have on the model’s prediction (e.g., vision and language biases). By performing controlled interventions of the images ($I$) and of the question ($Q$, $T$), we are able to quantify the causal effects of each factor.
In this section, we first introduce the causal graph of MLLM’s prediction on VQA problems. Then, we leverage the causal graph to elucidate the biases inherent in VQA, especially vision and language biases. Finally, we assess the causal effects of such biases on MLLMs’ prediction by performing controlled interventions Pearl (1995).
2.1 Problem Setup
We consider an entity-centric VQA problem, denoted by $M$, consisting of a question $Q$ and an image $I$. The image depicts a specific entity and the question is related to the entity. The question $Q:=(S,T)$ is composed of two distinct elements: the core semantic content ${S}$, which conveys the genuine intent of the question, and the textual surface form ${T}$, which is irrelevant to the question’s core meaning. The model’s final answer/prediction is represented by $A$. In this paper, we use lowercase letters to represent an instance of its corresponding capital variable.
2.2 Causal Graph of MLLMs’ Prediction
Inspired by the intuitive reasoning mechanisms observed in human cognition Zellers et al. (2019); Fei et al. (2022); Stolfo et al. (2022), we formulate the causal mechanisms underlying human problem-solving in a VQA problem $m$:
${s}=f_{\text{abstract }}({q}),\ \ g=f_{{s}}({i}),$ (1)
where they first abstract the core semantic meaning $s$ from the question $q$ by some cognitive process $f_{\text{abstract }}$, and then relate $s$ to the image $i$ to yield the final result $g$. We present these mechanisms in the green subgraph $G_{h}$ of Figure 2.
In contrast, the causal mechanisms of how a model might solve the same VQA problem $m$ are as follows:
$a=f_{\text{blackBox }}({i},{q}),\vspace{-5pt}$ (2)
where it is uncertain which aspect(s) of $q$ the model considers and how it interacts with the image $i$.
To conduct this analysis, we draw all possible causal mechanisms that might take place in the complete causal graph in Figure 2. Notable causal mechanisms include:
Vision Bias
The model may attend directly to the image $I$ via a causal path ${I}\rightarrow A$, leading to the emergence of vision bias, e.g., the focus on the image in Figure 1 (b).
Language Bias
The model may directly process the question $Q$ in two ways: by focusing on the core semantics $S$ via the causal path ${Q}\rightarrow S\rightarrow A$, or on the irrelevant part $T$ via the causal path ${Q}\rightarrow T\rightarrow A$. Both pathways lead to language bias, e.g., the focus on the question in Figure 1 (a).
Desired Causal Mechanisms
Intuitively, if the model learns the correct causal mechanisms to solve a VQA problem (i.e., the green subgraph $G_{h}$ of Figure 2), it should understand how both the image and the question contribute to the ground-truth result ${G}$ (via ${I}\rightarrow G$ and ${S}\rightarrow G$). Consequently, the model’s predictions should exhibit sensitivity and robustness to changes in the ground truth, namely ${G}\rightarrow A$. No spurious correlations should directly affect $A$ without going through the mediator $G$. Based on the above analyses, we articulate the concept of model sensitivity and robustness on VQA problems:
  • Sensitivity, which assesses if the model adjusts its predictions properly when the correct answer changes, i.e., $A$ responds to changes in $G$.
  • Robustness, which assesses the undesired direct causal effect of unimodal biases, e.g., $I\rightarrow A$ and $T\rightarrow A$, where lower effects indicate greater robustness to input changes that do not alter the correct answer.
2.3 Causal Analysis of VQA Biases
After formulating the cognitively-inspired subgraph $\mathcal{G}_{h}$ and defining the path of unimodal biases, we are able to quantify the causal effects of each causal factor on another by performing controlled interventions Pearl (1995).
Causal Interventions
In the context of VQA, we adopt the following interventions to quantify the causal effects of images and questions on the model’s predictions:
  • 1).
    Direct intervention on the image $I$ by replacing an alternative image $I^{\prime}$.
  • 2).
    Partially controllable interventions on $Q$. The question $Q$ can be modified in two ways: (i) both $S$ and $T$ are altered, or (ii) $T$ is altered but $S$ remains unaffected.
Formulation of Causal Effects
Next, we explain how we can obtain the causal effect from the interventions. Let us consider an intervention $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{I,Q,T\}$ and the VQA problem $M=\{I,Q\}$. We denote the distribution before the intervention $\mathbb{P}(A\mid{I},{Q})$ as $P$ and the distribution after the intervention as $P^{\prime}$.
Following the distributional definition of causal effect by Pearl (1995), we quantify the effect of factor $X$ in our causal graph using a distance metric $\delta$ between the distributions $P$ and $P^{\prime}$. That is,
$\operatorname{CE}=\delta\left(P,P^{\prime}\right),$ (3)
where $\operatorname{CE}$ denotes the causal effect and can be further divided into the total causal effect ($\operatorname{TCE}$, i.e., the joint effect through all the directed causal paths from a variable to another), or the direct causal effect ($\operatorname{DCE}$, i.e., the effect of the directed causal path from a variable to another without any intermediate variables) Pearl (2022).
Following  Stolfo et al. (2022), we quantify the causal effect of factor $X$ on the model’s answer $A$ by assessing the change in the predicted result:
$\delta_{\mathrm{cp}}\left(P,P^{\prime}\right):=\mathbb{I}\left(a\neq a^{\prime}\right)$ (4)
where $a=\arg\max_{x}P(x),\text{ and }a^{\prime}=\arg\max_{x}P^{\prime}(x)$, $\mathbb{I}$ denotes an indicator of the “change answer” event.
Causal Effects of the Images
When intervening on the images $I$, we can obtain the size of causal effect of $I$ on $A$, namely
$\operatorname{TCE}({I}\text{ on }A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I})}\left[\delta\left(P,P^{\prime}\right)\right],$ (5)
[…]
Note that, this $\operatorname{TCE}$ consist of two different paths of how $I$ affects $A$, as shown in Figure 2:
  • The path ${I}\rightarrow G\rightarrow A$ represents the desired decision-making route we aim for the model to adopt, where it responds to alterations in the ground-truth answer.
  • The path ${I}\rightarrow A$ illustrates a spurious correlation that the model might have learned, where it depends on certain visual contexts potentially linked to their prevalence in the training corpus.
We can quantify the $\operatorname{DCE}$ of $I$ on $A$, namely the strength of the path ${I}\rightarrow A$, by maintaining $G$ constant during each intervention on $I$:
$\operatorname{DCE}({I}\rightarrow A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I}\mid G)}\left[\delta\left(P,P^{\prime}\right)\right],$ (6)
[…]
Causal Effects of the Questions
As for questions, by intervening on $Q$, we can compute the total causal effect of $Q$ on $A$:
$\operatorname{TCE}({Q}\text{ on }A):=\mathbb{E}_{{q}^{\prime}\sim\mathbb{P}({Q})}\left[\delta\left(P,P^{\prime}\right)\right],$ (7)
[…]
Controlling for the core semantic meaning $S$ will instead allow us to obtain the $\operatorname{DCE}$ of the textual surface form $T$ on $A$:
$\operatorname{DCE}({T}\rightarrow A):=\mathbb{E}_{{q}^{\prime}\sim\mathbb{P}({Q}\mid S)}\left[\delta\left(P,P^{\prime}\right)\right],$ (8)
[…]
Note that, since there is no mediator between $T$ and $A$, the $\operatorname{DCE}({T}\rightarrow A$) is also $\operatorname{TCE}$ of $T$ on $A$. Due to that it is generally infeasible to enumerate all possible perturbations of $T$, we could achieve practical results over a certain subset of $T$ by intervening on $Q$ without affecting $S$ Stolfo et al. (2022). Besides, in the context of the VQA problem, we cannot only intervene on $S$ without affecting the textual surface $T$. Nevertheless, the causal influence of $S$ on $A$ can be understood by comparing two quantities we already know, i.e., $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.
Overall, calculating $\operatorname{TCE}$ helps us assess a model’s sensitivity (response to changes in ground truth), while $\operatorname{DCE}$ evaluates its robustness (stability of predictions against spurious correlations with fixed ground truth).
3 Constructing the MORE Dataset
Figure 3: Our framework for generating data of MORE. We first prepare the image source and link the visual entity in a knowledge graph. Then, motivated by the visual and language bias analysis through the causal lens, we construct multiple-choice questions that require MLLMs to overcome unimodal biases and conduct multi-hop reasoning in a sampled subgraph. We also generate the causal (reasoning) rationale for each instance to provide interpretability.
Based on the causal analysis in Section 2, we construct a novel MORE dataset that necessitates MLLMs to transcend unimodal biases and thoroughly integrate information from both text and images to select the correct answer. To ensure quality, we employ various evaluations to analyze the quality of the generated data. The data generation process is illustrated in Figure 3.
3.1 Images and Knowledge Graph Collection
We begin with an existing VQA dataset, INFOSEEK Chen et al. (2023), which links entities depicted in images to information sourced from Wikipedia, requiring a VQA model to answer a question about the associated entity. For instance in Figure 3, the entity associated with the showcased vehicle is “Chrysler PT Cruiser”, with the question being “What equipment or engine is used by this vehicle to provide power?” In this example, terms such as “Chrysler PT Cruiser”, “this vehicle”, and the relevant region in the image all refer to the same entity.
Based on the images and corresponding entity information, we identify all $n$-order neighbors ($n\in\{1,2\}$) related to the associated entity in a knowledge graph (KG) - Wikidata5M Wang et al. (2021), which is built upon Wikipedia data.
3.2 Subgraph Sampling
Motivated by the causal analysis discussed in Section 2, we aim to construct multi-hop queries that require overcoming unimodal biases to be answered correctly. To this end, we first identify a subgraph of an entity and its $n$-order neighbors in the KG. We then filter paths in this subgraph that meet two criteria: 1) Uniqueness of Paths: the path from the associated entity to the selected neighbor is unique, and 2) Shared-Type Relations: they share a same-type relation pointing to a unique entity, and these two pointed entities are not the same. The first criterion guarantees the uniqueness of the correct answer, while the second introduces interference and challenges MLLM’s reasoning ability.
For instance, in the subgraph of Figure 3, “Fiat 500X” is a first-order neighbor of “Chrysler PT Cruiser”, uniquely linked by the “followed by” relation. These two entities are linked to different entities “The Chrysler Corporation” and “Fiat Automobiles S.p.A” through the same type of relations “brand”. Meeting these criteria, we form the multi-hop query: “Chrysler PT Cruiser $\stackrel{{\scriptstyle\text{followed by}}}{{\longrightarrow}}$iat 500X $\stackrel{{\scriptstyle\text{brand}}}{{\longrightarrow}}$ Fiat Automobiles S.p.A”, with “Chrysler PT Cruiser” as the start and “Fiat Automobiles S.p.A” as the answer. In this way, other entities of the filtered subgraph may serve as distractors to challenge MLLMs’ reasoning ability, which will be discussed in the next section.
3.3 Multiple-Choice Question Construction
In this subsection, we detail the process of constructing multiple-choice questions with four candidate options.
Question Generation
After obtaining the subgraph that meets the criteria, we generate questions using the entity-relation path in it. Specifically, to obtain fluent and coherent questions, we feed the path into an LLM to produce the target question text. Our inputs include the names of start entity and all relations within the path. We adopt the in-context learning (ICL) technique Brown et al. (2020) and provide several examples to LLMs. The prompt template is in Appendix A.1. Comparing various LLMs with adjusted instructions, we find ChatGPT generates the highest-quality multi-hop questions, hence choosing its results for subsequent human evaluation. Finally, to prevent information leakage, entity names in the question are replaced with “this <OBJECT_NAME>”.
Language Bias Option
As discussed in Section 2.2, language bias refers to the model overly attending to the information from questions through the undesired path ${Q}\rightarrow S\rightarrow A$ or ${Q}\rightarrow T\rightarrow A$. To simulate such a situation, we query an MLLM with the generated question in a text-only setup. To ensure that the final options are identical for all MLLMs, we uniformly use the answers obtained from querying GPT-4V. The output for the mentioned example in Figure 3 is “Tesla”. The prompt template we use can be seen in Appendix A.2.
Vision Bias Option
To explore vision bias along the path $I\rightarrow A$, we include the visually associated entity name (e.g., “Chrysler PT Cruiser”) as an option. This allows us to observe whether the model directly selects it upon encountering an option that aligns with the visual information.
Semantic Misleading Option
Moreover, we introduce a semantic misleading option, such as “The Chrysler Corporation”, to challenge the multi-hop reasoning of MLLMs in KG. This option refers to the entity that is pointed by the relation commonly owned by both the associated entity and its sampled neighbor. For example in Figure 3, upon encountering a question about “brand” and “Chrysler PT Cruiser”, MLLMs might simply output a direct answer (e.g., “The Chrysler Corporation”), ignoring other constraints in the question (e.g., “followed by”), hence struggling to choose the correct answer (e.g., “Fiat Automobiles S.p.A.”).
Ground Truth Option
Corresponding to the causal path via ${I}\rightarrow G$ and ${S}\rightarrow G$, this option is the end entity of the entity-relation path (e.g., “Fiat Automobiles S.p.A.”). Finally, we check and ensure that each option is distinct from the remaining three to eliminate any potential overlap in samples.
Figure 4: Question quality of MORE compared to other VQA datasets in terms of lexical diversity and fluency.
3.4 Causal Rationale Generation
Furthermore, the entity-relation paths can assist in generating the reasoning rationale, termed as causal rationale, tailored to the question at hand. In this context, we have employed a heuristic rule-based approach, starting from the associated entity and generating a causal rationale step by step until reaching the ground truth answer. Templates of generation can be found in the Appendix A.3. These generated causal rationales can be utilized to verify whether the reasoning process of MLLMs is correct, thereby providing interpretability. They can also be used to fine-tune specialized MLLMs to enhance their multi-hop reasoning capabilities, with experimental results available in Section 5.3. The causal rationales could further be polished and refined through LLMs such as ChatGPT, and we leave this for future work.
3.5 Dataset Statistics
Statistics of Different Hops
Using the aforementioned construction approach, we automatically generate training data from the train set of INFOSEEK, and development/test data from the validation set of INFOSEEK. Table 4 in Appendix shows the statistics of our curated dataset MORE. For the purpose of generating questions with depths of 2-hop and 3-hop, we set $n=1,2$, respectively. We do not include longer-hop queries here due to concerns that their complexity will render the generated questions ambiguous or difficult to comprehend.
Question Distribution
We categorize the generated questions into distinct types based on their starting n-grams in Figure 17 in the Appendix. The MORE dataset showcases an extensive lexical diversity in the questions generated.
3.6 Quality Analysis
Question Quality
To ensure the quality of the comprising datasets, we analyze the lexical diversity and the fluency of the generated questions. The considered baselines and metrics are detailed in Appendix C.1. From Figure  4, MORE shows superiority in terms of lexical diversity and fluency, even in comparison to datasets created through human annotator-based question generation.
Human Evaluation
We also conduct a human evaluation to validate and assess the quality of the generated questions. Results in Appendix C.2 show that 90.6% of the generated questions are classified as valid by the annotators, further demonstrating the quality of our datasets.
4 Evaluating MLLMs on MORE
Model LLM # Params MORE (Two-Hop, acc (%)) MORE (Three-Hop), acc (%) MORE (Overall, acc (%))
Open-ended Multi-choice Open-ended Multi-choice Open-ended Multi-choice
Random / / / 25.0 / 25.0 / 25.0
BLIP-2 OPT 6.7B 4.0 16.4 1.4 15.4 2.7 15.9
InstructBlip Vicuna 13B 3.0 17.0 1.6 16.2 2.3 16.6
mPLUG-Owl Llama 7B 4.0 12.4 8.2 11.4 6.1 11.9
LLaVA Llama 13B 8.0 20.8 6.8 13.6 7.4 17.5
GPT-4V - - 15.8 25.6 15.3 23.2 15.6 24.4
Gemini Pro - - 14.2 33.5 10.1 24.4 12.2 28.9
Table 2: MLLMs’ results on the test set of MORE. We report the VQA accuracy (%) under the open-ended generation and the multi-choice settings on two-hop, three-hop, and all data, respectively.“-” denotes not released information.
4.1 Experimental Setup
Datasets
We use all the test data of MORE for evaluation. We adopt two different settings: 1) Open-ended. Ask the MLLM to generate answers based on the input image and question, or 2) Multi-choice. Provide MLLM with four options and let it choose the correct answer from them. The latter setting has a random baseline (accuracy of 25%).
Baselines
We evaluate various leading MLLMs on our MORE dataset in a zero-shot fashion, including two limited-access MLLMs: GPT-4V OpenAI (2023) and Gemini Pro Team et al. (2023), and four open-source MLLMs: BLIP-2 (6.7B)  Li et al. (2023), InstructBLIP (13B)  Dai et al. (2023), mPLUG-Owl (7B) Ye et al. (2023), and LLaVA (v1.5, 13B) Liu et al. (2023a), detailed in Appendix D.1. As for evaluation metrics, we adopt the VQA accuracy Antol et al. (2015) for all models for a fair comparison. We also quantify the causal effects of images and questions on models’ predictions in Section 4.3.
4.2 Evaluation Results
We present the results of MLLMs separately on two-hop, three-hop, and all data of MORE in Table 2. Examples of model prediction can be found in Appendix H. We observe that: 1) The performance of all baselines on MORE is poor (e.g., under the “Multi-choice” setting, only Gemini Pro exceeded the random baseline at 28.9% accuracy), indicating MLLMs’ vulnerability to language and vision biases. 2) There is still a gap between open-source and limited-access models on MORE, especially under the “Open-ended” setting. 3) Most models perform better on two-hop data than on three-hop data (Gemini Pro is particularly outstanding on two-hop data, achieving 33.5% accuracy), suggesting that MLLMs’ reasoning capabilities are challenged when the problems become more complex. 4) GPT-4V performs best under the “Open-ended” setting, but falls short under the “Multi-choice” setting versus Gemini Pro, possibly due to that we use homologous ChatGPT-generated distractors when constructing the language bias option in Section 3.3, which poses a greater challenge to GPT-4V’s judgment. This point is also verified in the following analyses in Section 4.3.
4.3 Causal Analysis of VQA Biases
In this subsection, we analyze the performance of MLLMs through a causal lens.
Option Distribution
Figrue 5 shows the distribution of options under the “Multi-choice” setting across various MLLMs. We observe that: 1) BLIP2 and GPT-4V often incorrectly choose options indicating language bias, aligning with our prior analysis on GPT-4V. 2) The proportion of either language or vision bias exceeded 40% across all the models, demonstrating the significant impact of unimodal biases on their predictions. 3) To some extent, models’ selection of semantically misleading options indicates some ability to combine visual and textual information, though not fully grasping the problem. This highlights the challenge our MORE dataset poses to current MLLMs. Please note that discrepancies may exist between the proportions of ground truth options presented here and the accuracy values reported in Table 2, as some models’ (e.g., mPLUG-Owl) outputs may not align with the provided options, affecting the count of valid answers.
Causal Effects of Images and Questions
Figure 5: Option distribution of MLLMs.
Figure 6: Comparison of direct and total effects of image and question on prediction for MLLMs.
To further analyze the impact of vision bias and language bias on the model’s prediction, we assess the causal effects based on the definitions provided in Section 2.3. Specifically, we randomly select 100 samples for intervention and then measure the average of the effects over all instances to calculate $\operatorname{TCE}$ (corresponds to model’s sensitivity) and $\operatorname{DCE}$ (corresponds to models’ robustness). Overall, a higher $\operatorname{TCE}$ is desirable, indicating better sensitivity, while a lower $\operatorname{DCE}$ indicates better robustness. Implementation details of the intervention can be found in Appendix E.
From Figure 6: 1) Current MLLMs exhibit high sensitivity (high $\operatorname{TCE}$), a possible reason is that instruction tuning makes models sensitive to variations in input Stolfo et al. (2022). 2) However, the robustness is relatively low (high $\operatorname{DCE}$), showing predictions change with variations in input even when the ground truth is fixed, suggesting the reliance on spurious paths over genuine causal features. 3) The relationship between MLLMs’ utilization of biases and their performance is intricate, with GPT-4V performing well even amid strong language and vision biases.
5 Mitigating VQA Biases for MLLMs
Model / Acc (%) OK-VQA INFOSEEK MORE
Open-ended. Multi-choice.
GPT-4V 70.4 38.0 15.6 24.4
Gemini Pro 64.9 37.8 12.2 28.9
LLaVA 72.8 16.0 7.4 17.5
GPT-4V (w. the DeVA framework) 74.8 50.6 24.2 30.3
Gemini Pro (w. the DeVA framework) 68.3 50.3 19.5 32.1
LLaVA (trained on INFOSEEK) 72.1 19.7 3.1 13.9
LLaVA (trained on MORE) 73.6 20.2 16.4 28.9
Table 3: Results comparison after employing our strategies on OK-VQA, INFOSEEK, and MORE.
In this section, we explore two strategies to mitigate the unimodal biases of MLLMs.
5.1 Experimental Setup
Dataset
We conduct experiments on three knowledge-based VQA datasets OK-VQA Marino et al. (2019), INFOSEEK Chen et al. (2023), and our curated MORE, where relying solely on image content proved insufficient for answering questions. We randomly select 1,000 instances from the validation set of INFOSEEK and OK-VQA, respectively, and report results on the test data of MORE, to ensure a consistent number of samples for evaluation. Since INFOSEEK and OK-VQA do not provide multiple options, we only report the results of them under the “Open-ended” settings.
Baselines
We focus on three models that have shown superior performance in Table 2: LLaVA, GPT-4V, and Gemini Pro. For limited-access models GPT-4V and Gemini Pro, we propose a prompting-based Decompose-Verify-Answer (DeVA) framework. For the open-source model LLaVA, we conduct fine-tuning on various datasets and compare the results. We do not apply the DeVA framework to LLaVA based on findings from our pilot study that LLaVA’s performance in comprehending and responding to our instructions is not on par with that of GPT-4V and Gemini Pro.
5.2 A Decompose-Verify-Answer Framework
For a given instance, DeVA begins with a question decomposer to break down the complex question into several easier ones step-by-step, so as to explicitly avoid the model simply taking a spurious path to give the answer. For each subquestion, after providing an answer, DeVA employs a verifier to confirm the accuracy of the answer. The verifier strategizes the utilization of external tools and investigates their outputs, thereby acquiring the necessary context or information needed to provide a precise answer, such as image retrieval and text retrieval. This iterative process of answering and verifying will continue until all subquestions have been resolved. By incorporating external knowledge, the final verified output comprehensively integrated the information of the whole reasoning process to give a correct answer. A detailed illustration of DeVA and the prompt template is in Appendix F.
Results
From Table 3, we can see that: 1) the DeVA framework significantly enhances the performance of GPT-4V and Gemini Pro across all three datasets. 2) Notably, the improvement on the INFOSEEK dataset is more pronounced than on MORE. This is because INFOSEEK is a dataset focused on information-seeking, and once we provide the models with relevant external knowledge, they can quickly extract the necessary information and output a correct answer. However, the limited results on MORE suggest that even when models are allowed access to additional information, they may still be impeded by language and vision biases. Therefore, how to disentangle and address this issue remains a challenging problem, which we leave for future work.
5.3 Fine-tuning LLaVA
We fine-tune LLaVA using the LoRA Hu et al. (2022) technique on the training sets of INFOSEEK and MORE, respectively. For MORE, we incorporate the generated causal rationale into the instruction of some samples, so as to enhance the models’ reasoning ability. Other implementation details are in Appendix G. From Table 3, we can see: 1) After fine-tuning, LLaVA shows a significant improvement in performance on OK-VQA, INFOSEEK, and MORE, particularly LLaVA (trained on MORE) outperforms the vanilla GPT-4V on OK-VQA and MORE. 2) While LLaVA (trained on INFOSEEK) does improve its performance on INFOSEEK, it significantly reduces its effectiveness on MORE. In contrast, LLaVA (trained on MORE) demonstrates better generalizability, confirming the effectiveness of incorporating rationales into instructions.
6 Related Work
Multimodal Large Language Models
Recent advancements in LLMs have facilitated the development of powerful MLLMs. Pioneer work has demonstrated their impressive capability on multimodal tasks OpenAI (2023); Team et al. (2023); Liu et al. (2023a). Despite the substantial success, the current evaluation of MLLMs mainly focuses on basic visual tasks Liu et al. (2023b); Fu et al. (2023); Lu et al. (2024), and their reasoning ability has not been investigated in detail.
Knowledge-based VQA Datasets
Several previous works have studied knowledge-based VQA problems Wang et al. (2017); Marino et al. (2019); Chen et al. (2023). While these datasets do require knowledge to answer questions, they have the following limitations: 1) They often focus solely on image-related information without requiring the model to perform multi-hop reasoning. 2) The answer types are typically open-ended and do not provide reasoning rationales, making it difficult to evaluate the model’s output. 3) They cannot assess the degree of reliance on language or visual biases by models. As initially introduced in Section 1, Table 1 provides a comparative summary of various datasets alongside our MORE dataset, underscoring its enhanced comprehensiveness.
Language and Vision Biases in VQA
Studies have found that some VQA models tend to rely on statistical priors present in the training data, rather than truly understanding the content of images Agrawal et al. (2018). This issue manifests in two primary forms: language bias and vision bias. The former emerges from strong correlations between specific questions and their answers Abbasnejad et al. (2020); Zhu et al. (2020), and the latter stems from the frequent co-occurrence of certain textual and visual elements in the training dataset Si et al. (2022); Gupta et al. (2022). Recent strategies to mitigate such biases mainly focus on employing data augmentation techniques Niu et al. (2021). Although the concerns are similar, all the aforementioned methods focus on the statistical bias under the pre-train and fine-tune paradigm, while we focus on the reasoning ability of MLLMs.
7 Conclusion
This paper presents a comprehensive approach to quantifying and mitigating the unimodal biases in MLLMs. Through our causal inference framework, we provide an in-depth analysis to assess the causal effects of such biases on the model’s prediction in VQA problems. The introduced MORE dataset challenges MLLMs to engage in multi-hop reasoning and to overcome language and vision biases, thereby pushing the boundaries of their reasoning capabilities. Our proposed solutions, including the DeVA framework and the fine-tuned LLaVA models, demonstrate significant potential in enhancing the reasoning abilities of MLLMs.
8 Limitations
Our current generation of rationales is based on heuristic rules. Experiments in Section 5.3 have demonstrated the effectiveness of incorporating rationales into instructions. Therefore, we believe that refining and polishing these rationales with an LLM (e.g., ChatGPT) could be beneficial. Besides, the Wikidata5M dataset we employed was released in 2021, and some information in the knowledge graph may be outdated. Although we have made efforts to manually verify the test set and try to ensure it does not contain incorrect information, it is still inevitable that errors may occur within the extensive training data.
Appendix A Prompt Templates
A.1 Question Generation
Figure 7: Prompt template of multi-hop question generation.
We present the prompt template for generating language bias options in Section 3.3 in Figure 7.
A.2 Language Bias Option Generation
Figure 8: Prompt template of language bias option generation.
We present the prompt template for generating language bias options in Section 3.3 in Figure 8.
A.3 Training Instance with Causal Rationale
Figure 9: Prompt template of the training instance in MORE.
Figure 10: Human evaluation results of MORE.
An example of the training instance in MORE is shown in Figure 9.
Appendix B Question Distribution
In Figure 17, we categorize the generated questions into distinct types, based on their starting n-grams. The dataset MORE showcases an extensive lexical diversity in the questions generated. This diversity is evidenced by variations in the introductory interrogative words (e.g., “what”, “who”, “where”, etc.), exemplified by phrases like “What is the…”, “In which country…”, and more. Such lexical richness is crucial for mitigating the vulnerability of MLLMs to linguistic variations.
Appendix C Quality Analysis Details
C.1 Question Quality
To ensure the quality of the comprising datasets, we analyze the lexical diversity and the fluency of the generated questions, which are useful for conducting a robust evaluation using questions that are linguistically diverse and coherent.
Baselines
We select extensive VQA datasets for comparison, including Visual7W Zhu et al. (2016), VQA (v2) Goyal et al. (2017), FVQA Wang et al. (2017), OK-VQA Marino et al. (2019), S3VQA Jain et al. (2021), A-OKVQA Schwenk et al. (2022), and INFOSEEK Chen et al. (2023) (contains both automated generation version and human-annotated version).
Evaluation Metrics
For lexical diversity, we utilize three metrics that are not dependent on length: moving average type-token ratio (MATTR) Covington and McFall (2010), measure of textual lexical diversity (MTLD) McCarthy (2005), and hypergeometric distribution diversity (HDD) McCarthy and Jarvis (2010). We average these three metrics for a unified assessment and employ the Lexical-Richness package Shen (2022) (version 0.5.03) for calculation. For fluency, we employ a pre-trained language model GPT2-large Radford et al. (2019) with 774M parameters to compute the perplexity of the questions, which is often used as a measure by previous work Wang et al. (2019); Cahyawijaya et al. (2021).
C.2 Human Evaluation
We conduct a human evaluation of 100 questions randomly chosen from the MORE dataset to validate and assess the quality of the generated questions. This evaluation is carried out by three human annotators, who are provided with detailed guidelines and illustrative examples before starting the evaluation process. For each question, given the visual context and ground truth answer, the annotators are asked to determine whether: 1) the question is valid, 2) the question has a possible answer alternative, 3) the question does not match the answer, or 4) the question is unclear or confusing. For each sample, we ask three human annotators to judge each generated question into one of the four options.
The outcomes of this human evaluation are summarized in Figure 10. The results are encouraging, with 90.6% of the generated questions being classified as valid by the annotators, further demonstrating the quality of our datasets.
Appendix D Baselines and Datasets Details
D.1 Baselines
For open-source MLLMs, we consider the following baselines:
1) BLIP2 Li et al. (2023), which utilizes a scalable multimodal pre-training method to enable LLMs to understand images. We employ its BLIP2-OPT Zhang et al. (2022)-6.7B variant.
2) InstructBLIP Dai et al. (2023), an instruction-tuned version of BLIP-2 on various tasks including VQA. We employ its InstructBLIP-Vicuna Chiang et al. (2023)-13B variant.
3) mPLUG-Owl Ye et al. (2023), which proposes a new two-stage training method for aligning images and text. We employ its mPLUG-Owl-Llama Touvron et al. (2023a)-7B variant.
4) LLaVA Liu et al. (2023a), which translates images into texts of captions and bounding boxes, and prompts GPT-4 to generate a multimodal instruct-tuning dataset in the context of seed examples. We employ its LLaVA-Llama Touvron et al. (2023a)-13B variant.
D.2 Datasets
OK-VQA Marino et al. (2019) is a VQA dataset that necessitates external knowledge beyond the visible content in images. It is crafted by crowdsourced participants who form complex questions based on extensive information available on Wikipedia.
INFOSEEK Chen et al. (2023) is another VQA dataset aimed at information-seeking questions that require more than just commonsense knowledge. It combines human-annotated questions with visual entity recognition datasets and Wikidata to generate complex question-answer pairs.
Appendix E Implementation Details of Intervention
In this section, we introduce the implementation details of interventions on images and questions.
For images, in the analysis of $\operatorname{TCE}$, we maintain a constant question while replacing the image with another from the dataset that corresponds to the same question but features different entities, ensuring a change in the ground truth post-intervention. In the analysis of $\operatorname{DCE}$, the question remains unchanged, but the image is replaced with one of the same entity from an alternative source, specifically using Google Lens’s API***Web interface available at https://images.google.com.) to keep the ground truth consistent after the intervention.
For questions, in the analysis of $\operatorname{TCE}$, the image stays the same while the question is altered to a different one from the dataset that relates to the same entity, ensuring a ground truth shift post-intervention. In the analysis of $\operatorname{DCE}$, we similarly keep the image constant and let ChatGPT rephrase the question, altering its textual form but preserving its semantic meaning (verified through manual evaluation), thereby maintaining an unchanged ground truth after the intervention.
Appendix F The DeVA Framework
In this section, we first present the overall framework, then we will go over each part of it in detail.
F.1 Overall Framework
Given a question $Q$ and an image $I$, the VQA task demands the system to return an output $A$ that concisely answers the question. As shown in Figure 11, we first initialize a question decomposer $D$ to analyze $Q$ and break it down into manageable subquestions. Then, we employ a verifier $V$ to confirm the accuracy of the original answer to each decomposed subquestion. The generation verifier typically involves active information-seeking and answer-verification, which acquires the necessary context or information needed to investigate and revise the answers. This includes two optional operations: image retrieval to seek images similar to $I$ and determine their titles, or text retrieval with a specific query to fetch pertinent documents and summarize their content. This iterative process of answering and verifying will continue until we resolve each subquestion. Finally, the verified answer $A$ is output following the aforementioned reasoning process and retrieved information.
F.2 Question Decomposer
For a given VQA problem, MLLMs often simply exploit a spurious path to make predictions due to the inherent language and vision biases. In order to alleviate this issue, motivated by Chain-of-Thought reasoning Wei et al. (2022), we encourage MLLMs to decompose the question $Q$ before outputting the answer, so as to gradually solve a complex question that requires multi-hop reasoning. As shown in Figure 11, for the question “Which country hosted the next World Cup after this venue?”, our decomposer breaks it down into two subquestions :
  • 1.
    What this venue is?
  • 2.
    Which country hosted the next World Cup after Allianz Arena?
Such decomposition will explicitly constrain the model to comprehend and extract the truth semantics of the question, thus avoiding simply exploring a spurious path to give the answer.
F.3 Verifier
Some works have found that vision illusion and language hallucination may appear in the process of MLLMs’ response generation Guan et al. (2023); Yang et al. (2023). To alleviate this issue, we adopt the retrieval-augmented generation approaches Khandelwal et al. (2020); Chen et al. (2022). Specifically, we consider two different retrieval ways for the verifier to choose during each verification step: image retrieval and text retrieval.
Image Retrieval
Although our framework is applicable to any image retrieval method, in this paper, we mainly utilize Google Image Search to obtain a broad range of information related to the image as provided by Google Lens API. This information encompasses various details, such as knowledge graph entities and captions of analogous or identical images. The availability of these details can vary based on the image input provided to Google Image Search. Then, the verifier gleans relevant information from captions associated with visually similar images, so as to verify the original answer and conduct the next round of reasoning.
Text Retrieval
Similarly, our framework is applicable to any text retrieval method, we explore a simple, off-the-shelf dense retriever for Wikipedia, GTR Karpukhin et al. (2020), as our text retriever. First, the verifier constructs a query to perform text retrieval according to the currently generated context, and then the query is input into a GTR model to get related document titles and contents. Finally, the verifier will fetch pertinent documents and summarize their content to verify the immediate answer.
Figure 11: An overview of our proposed Decompose-Verify-Answer (DeVA) Framework.
Figure 12: Prompt template of DeVA framework.
F.4 Final Verified Answer
Finally, the improved response that takes verification into account is generated. This is executed by a final prompt where the context takes into account all of the previous reasoning steps, the baseline response, and the verification question-answer pairs, so that the corrections can take place. The prompt template is shown in Figure 12.
Appendix G Implementation Details of Fine-tuning LLaVA
We adopt LLaVA-13B-v1.5 as the base model and employ the LoRA Hu et al. (2022) technique. During pre-training, only LoRA parameters are optimized. We set the rank of LoRA modules to 128. Our model is optimized with a learning rate of 2e-4 and a linear warm-up for the first 3% steps. The batch size is 16 and the number of epochs is 1. All the LoRA parameters are trained on 8 NVIDIA A100 GPUs with 80GB memory. We fine-tune LLaVA on the training sets of INFOSEEK and MORE, respectively. For MORE, we incorporate the generated rationale into the instructions of some samples, so as to enhance the models’ reasoning ability. The template of a training instance can be found in Figure 9.
Appendix H Case Study
We conduct a case study on the development set of MORE in Figure 13$\sim$16, including both the “Open-ended” and “Multi-Choice” settings.
Figure 13: Two-hop case in the dev set of MORE. The wrong answers are marked in red and the correct is in green.
Figure 14: Two-hop case (multi-choice setting) in the dev set of MORE. The wrong answers are marked in red and the correct is in green.
Figure 15: Three-hop case in the dev set of MORE. The wrong answers are marked in red.
Figure 16: Three-hop case (multi-choice setting) in the dev set of MORE. The wrong answers are marked in red and the correct is in green.
Figure 17: Question prefix distribution in MORE. The arc length is proportional to the number of questions containing the word.