### v4

# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

Abstract
Recent advancements in Large Language Models (LLMs) have facilitated the development of Multimodal LLMs (MLLMs). Despite their impressive capabilities, MLLMs often suffer from over-reliance on unimodal biases (e.g., language bias and vision bias), leading to incorrect answers or hallucinations in complex multimodal tasks. To investigate this issue, we propose a causal framework to interpret the biases in Visual Question Answering (VQA) problems. Within this framework, we conduct an in-depth causal analysis to assess the causal effect of these biases on MLLM predictions. Based on the analysis, we introduce 1) a novel MORE dataset with 12,000 challenging VQA instances requiring multi-hop reasoning and overcoming unimodal biases. 2) a causality-enhanced agent framework CAVE that guides models to comprehensively integrate information from different modalities and mitigate biases. Our experiments show that MLLMs perform poorly on MORE, indicating strong unimodal biases and limited semantic understanding. However, when integrated with our CAVE, promising improvements in reasoning and bias mitigation can be seen. These findings provide important insights for the development of more robust MLLMs and contribute to the broader goal of advancing multimodal AI systems capable of deeper understanding and reasoning. Our project page is at https://github.com/OpenCausaLab/MORE.
Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective
Meiqi Chen1,2111This work was done during her internship at Shanghai AI Laboratory., Yixin Cao3, Yan Zhang1,$2$Corresponding author., Chaochao Lu4222Corresponding author. 1State Key Laboratory of General Artificial Intelligence, Peking University, Beijing, China 2School of Intelligence Science and Technology, Peking University 3School of Computer Science, Fudan University 4Shanghai Artificial Intelligence Laboratory meiqichen@stu.pku.edu.cn, caoyixin2011@gmail.com, zhyzhy001@pku.edu.cn, luchaochao@pjlab.org.cn
1 Introduction
Following the success of Large Language Models (LLMs) Ouyang et al. (2022); Touvron et al. (2023b), Multimodal LLMs (MLLMs) OpenAI (2023); Team et al. (2023) have been proposed for various vision-language tasks Fu et al. (2023); Liu et al. (2023b). Despite their promising results, it remains unclear if they truly understand images and text in the context of multimodal reasoning.
Figure 1: Examples of over-reliance on unimodal biases. MLLMs (e.g., LLaVA) erroneously generate answers due to language bias (indicated by the underlined text below the left image) and vision bias (the right image).
As shown in the knowledge-based Visual Question Answering (VQA) problems of Figure 1, when prompted with “Which country is hosting the next World Cup after this venue?” MLLMs such as GPT-4V OpenAI (2023) and LLaVA Liu et al. (2023a) may capture the language bias of “the next World Cup” and think that the next World Cup will be “the 2022 FIFA World Cup held in Qatar” (which is also outdated knowledge), while ignoring the exact venue presented in the image. Similarly, when presented with an image of “The Shard” in London, MLLM directly identifies “The representative building is The Shard” influenced by vision bias, overlooking the specific constraint “in Berlin” mentioned in the question. These inherent issues pose significant challenges to the reasoning capabilities of MLLMs, particularly when faced with more complex questions.
To investigate the issue of MLLMs’ Over-REliance (MORE) on such unimodal biases, we propose a causal framework to interpret and quantify language and vision biases. We begin by defining a causal graph of MLLM’s prediction on VQA problems, built on key causal factors like images and questions. Then, we identify a set of interventions in the context of VQA problems, thereby ascertaining the causal effect of unimodal biases on MLLM predictions via $do$-calculus Pearl (1995). This allows us to evaluate the sensitivity and robustness of MLLMs against unimodal biases.
Datasets
Knowledge-
based
Multi-hop
Reasoning
Answer Type
Unimodal Biases
Evaluation
Rationale # Size
Visual7W  Zhu et al. (2016) Open-ended 327.9K
VQA (v2) Goyal et al. (2017) Open-ended 1.1M
FVQA Wang et al. (2017) Open-ended 5.8K
OKVQA Marino et al. (2019) Open-ended 14K
S3VQA Jain et al. (2021) Open-ended 7.5K
A-OKVQA Schwenk et al. (2022) Multi-choice 23.7K
INFOSEEK Chen et al. (2023) Open-ended 1.4M
MORE (Ours) Multi-choice 12K
Table 1: Comparison of MORE with other VQA datasets, highlighting its incorporation of external knowledge, multi-hop reasoning, unimodal bias evaluation, and rationale for interpretability.
Based on the above causal analysis, we curate a novel dataset termed MORE, comprising 12,000 VQA instances. This dataset advances existing VQA datasets by introducing a dedicated evaluation of unimodal biases. We adopt a Multiple Choice Question (MCQ) format to facilitate the evaluation, where each instance consists of an image, a question, and four candidate options. The image is sourced from an existing VQA dataset Chen et al. (2023). For question and option curation, we incorporate a knowledge graph (KG) Wang et al. (2021), allowing us to simulate MLLMs to navigate potential paths within the causal graph. Specifically, the options consist of one correct answer, and three distractors targeting language bias, vision bias, and multi-hop reasoning, respectively. We also provide the reasoning path, designated as causal rationale, in the KG for each instance, offering interpretability for evaluation. As summarized in Table 1, compared to existing VQA datasets, MORE features better comprehensiveness.
Furthermore, motivated by the causal analysis, we propose CAVE, a causality-enhanced method to mitigate unimodal biases in MLLMs. CAVE encompasses a diverse set of workflows, including question decomposition, causality-based enhanced self-reflection, external knowledge retrieval, and answer verification. This framework guides models in explicitly and comprehensively integrating information from multiple modalities while helping to prevent biased decision-making and the selection of incorrect shortcuts.
Through extensive experiments on MORE with several leading MLLMs, we observe that: 1) most MLLMs perform much poorly on MORE, showing strong reliance on unimodal biases and low robustness to disturbances; 2) MLLMs still struggle to achieve precise semantic understanding in multimodal reasoning; 3) With our propose method CAVE, we can mitigate the unimodal bias in MLLMs, yet it still falls short of the ideal. This indicates that addressing the unimodal bias in MLLMs is a highly challenging issue that merits further exploration. Overall, our main contributions are as follows:
  • We propose a causal framework to interpret and quantify the biases in VQA problems.
  • We construct a new dataset, MORE, which requires multi-hop reasoning and overcoming biases, demonstrating superior comprehensiveness to existing VQA datasets.
  • We conduct extensive experiments on MORE and propose a causality-enhanced method CAVE to mitigate the unimodal biases, providing insights for future work.
2 A Causal Framework
Figure 2: Causal graph of MLLM’s Prediction on VQA problems. We use the green subgraph $G_{h}$ to represent the desired causal mechanisms and compare it with the undesired effects of unimodal biases. We quantify the causal effects of each factor by performing controlled interventions of the images ($I,E,C$) and of the questions ($Q,T$).
Inspired by Stolfo et al. (2023), we introduce a causal graph for MLLM predictions on VQA problems, highlighting language and vision biases and assessing their causal effects via controlled interventions Pearl (1995).
2.1 Problem Setup
We consider an entity-centric VQA problem, $M$, involving a question $Q:=(S,T)$ and an image $I:=(E,C)$. Here, $S$ is the core semantic content of $Q$, and $T$ is the textual form unrelated to $Q$’s main meaning. $I$ includes the primary entity $E$ and surrounding visual context $C$. The model outputs an answer, $A$. Notation: lowercase letters denote instances of corresponding uppercase variables.
2.2 Causal Graph of MLLM Predictions
Inspired by human cognitive processes Zellers et al. (2019); Stolfo et al. (2023), we formulate the causal mechanisms in human problem-solving for a VQA problem $m$: ${s}=f_{c_{1}}({q}),\ {e}=f_{c_{2}}({i}),\ g=f_{c_{3}}({s,e})$. This involves decoding the question $q$ to extract its semantic meaning $s$ through cognitive process $f_{c_{1}}$, and identifying the core entity $e$ from the image $i$ through $f_{c_{2}}$. These are combined by $f_{c_{3}}$ to produce the result $g$, as shown in the green subgraph $G_{h}$ of Figure 2. In contrast, the model’s approach to the same VQA problem employs $a=f_{b}({q},{i})$, a black-box function where it is unclear how question and image inputs are utilized and interact to form the prediction $a$. For deeper analysis, we illustrate possible causal mechanisms in Figure 2. Notable mechanisms include:
Language Bias
The model may directly process the question $Q$ in two ways: by focusing on the core semantics $S$ via ${Q}\rightarrow S\rightarrow A$, or on the irrelevant part $T$ via ${Q}\rightarrow T\rightarrow A$. Both pathways lead to language bias, e.g., the focus on “the next World Cup ” in Figure 1 (a).
Vision Bias
The model may attend directly to the entity $E$ of the image $I$ via $I\rightarrow{E}\rightarrow A$, or to the irrelevant part $C$ via $I\rightarrow{C}\rightarrow A$. Both pathways lead to the emergence of vision bias, e.g., the focus on the “The Shard” entity in Figure 1 (b).
Desired Causal Mechanisms
Correct reasoning in VQA problems hinges on understanding causal mechanisms, as depicted in Figure 2. The subgraph $G_{h}$ represents comprehension of how image and question jointly affect the ground-truth result ${G}$, through ${E}\rightarrow G$ and ${S}\rightarrow G$. This understanding should lead to model predictions that are sensitive and robust to changes in ${G}$, namely ${G}\rightarrow A$, with no spurious effects on $A$ unless it passes the mediator $G$. Model performance could therefore be evaluated by its: 1) Sensitivity, which assesses how well the model adapts to changes in the correct answer, i.e., $A$ responds to changes in $G$. 2) Robustness, which measures resistance to unimodal biases, e.g., $C\rightarrow A$ and $T\rightarrow A$, where minimal bias effects imply higher robustness to input changes that do not alter the ground-truth answer.
2.3 Causal Analysis of VQA Biases
We adopt the controlled interventions as outlined by Pearl (1995) to quantify the causal effects of questions and images on model predictions.
Causal Interventions for VQA
1) Interventions on $Q$. The question $Q$ can be modified in two ways: (i) altering both $S$ and $T$ , or (ii) altering $T$ but keeping $S$ unaffected. 2) Interventions on $I$. The image $I$ can be replaced with an alternative image $I^{\prime}$ in three ways: (i) altering both $E$ and $C$ , or (ii) altering $C$ but keeping $E$ unaffected, or (iii) altering $E$ but keeping $C$ unaffected. Note that we do not solely alter $S$ within $Q$, because it is not feasible to intervene on the core semantics $S$ of a question without affecting the surface text $T$ of it.
Formulation of Causal Effects
We assess the causal effects using $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{Q,T,I,C,E\}$ is a factor in the VQA problem $M$. The pre-intervention probability distribution $\mathbb{P}(A\mid I,Q)$ is denoted as $P$, and the post-intervention distribution as $P^{\prime}$. Following Pearl (1995), the causal effect is evaluated using a distance metric $\delta$, as $\operatorname{CE}=\delta(P,P^{\prime})$, where $\operatorname{CE}$ denotes the causal effect. It can refer to 1) the total causal effect ($\operatorname{TCE}$), signifying the joint effect across all causal paths from one variable to another; or 2) the direct causal effect ($\operatorname{DCE}$), indicating the effect of the directed causal path devoid of intermediary variables Pearl (2022). Following Stolfo et al. (2023), we assess the causal effect of factor $X$ on the model’s answer $A$ by comparing the change in predicted results, $\delta_{\mathrm{cp}}(P,P^{\prime}):=\mathbb{I}(a\neq a^{\prime})$, where $a=\arg\max_{x}P(x)$ and $a^{\prime}=\arg\max_{x}P^{\prime}(x)$, with $\mathbb{I}$ indicating a change in the answer.
Causal Effects of Questions
We assess $\operatorname{TCE}$ of a question $Q$ on an answer $A$ by intervening on $Q$:
$\operatorname{TCE}({Q}\text{ on }A):=\mathbb{E}_{{q}^{\prime}\sim\mathbb{P}({Q})}\left[\delta\left(P,P^{\prime}\right)\right],$ (1)
[…]
This $\operatorname{TCE}$ contains two different types of paths that show how $Q$ affects $A$, as illustrated in Figure 2: 1) The intended decision-making pathway: ${Q}\rightarrow S\rightarrow G\rightarrow A$, responding to changes in the ground truth. 2) Potential spurious correlations, e.g., ${Q}\rightarrow T\rightarrow A$, where the model may depend on certain linguistic patterns from training data.
Maintaining the core semantics $S$ constant, we can assess $\operatorname{DCE}$ of the textual surface $T$ on $A$:
$\operatorname{DCE}({T}\rightarrow A):=\mathbb{E}_{{q}^{\prime}\sim\mathbb{P}({Q}\mid S)}\left[\delta\left(P,P^{\prime}\right)\right],$ (2)
[…]
As discussed in “Causal Interventions for VQA”, solely intervening on $S$ without altering $T$ is impractical. However, Understanding $S$’s causal impact on $A$ is feasible by comparing two known quantities: $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.
Causal Effects of Images
The causal structure of images mirrors that of questions, as shown in Figure 2. Thus, we can derive $\operatorname{TCE}$ of image $I$ on the answer as follows:
$\operatorname{TCE}({I}\text{ on }A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I})}\left[\delta\left(P,P^{\prime}\right)\right],$ (3)
[…]
Likewise, maintaining $E$ constant during each intervention on $I$ allows us to quantify the $\operatorname{DCE}$ of the irrelevant visual context $C$ on $A$:
$\operatorname{DCE}({C}\rightarrow A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I}\mid E)}\left[\delta\left(P,P^{\prime}\right)\right],$ (4)
[…]
Maintaining $C$ and $G$ constant during each intervention on $I$ instead allows us to quantify the $\operatorname{DCE}$ of the core entity $E$ on $A$:
$\operatorname{DCE}({E}\rightarrow A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I}\mid C,G)}\left[\delta\left(P,P^{\prime}\right)\right],$ (5)
[…]
Overall, calculating $\operatorname{TCE}$ helps us assess models’ sensitivity (response to changes in ground truth), while $\operatorname{DCE}$ evaluates its robustness (stability of predictions against spurious correlations).
3 the MORE Dataset
Figure 3: Our framework for generating data of MORE. We first prepare the image source and link the visual entity in a knowledge graph. Then, motivated by the visual and language bias analysis through the causal lens, we construct multiple-choice questions that require MLLMs to overcome unimodal biases and conduct multi-hop reasoning in a sampled subgraph. We also generate the causal (reasoning) rationale for each instance to provide interpretability.
From Section 3.1 to 3.3, we construct the MORE dataset that exploits the unimodal biases defined in Section 2.2 and requires multi-hop reasoning. In Section 3.4, we discuss how to quantify the causal effects of images and questions (defined in Section 2.3) using our constructed dataset.
3.1 Preparatory Steps
Images and Knowledge Graph Collection
We begin with an existing VQA dataset, INFOSEEK Chen et al. (2023), which links image entities to Wikipedia information, requiring a VQA model to answer related questions. For instance in Figure 3, the entity “Chrysler PT Cruiser” prompts the question, “What equipment or engine is used by this vehicle to provide power?” Here, terms like “Chrysler PT Cruiser” and “this vehicle” all refer to the same entity in the image. We then identify all $n$-order neighbors ($n\in\{1,2\}$) of the associated entity within a knowledge graph (KG), Wikidata5M Wang et al. (2021), which is built upon Wikipedia data.
Subgraph Sampling
[…]
3.2 Multiple-Choice Question Construction
We detail the multiple-choice questions construction with four options, guided by the unimodal bias definition in Section 2.2.
Question Generation
We generate questions by analyzing entity-relation paths within a specified subgraph, converting these paths into fluent, coherent queries with an LLM. We utilize the in-context learning (ICL) technique Brown et al. (2020) and a standardized prompt (shown in Appendix A.1) to ensure quality. Among various LLMs tested, ChatGPT was selected for its superior multi-hop question generation. To avoid information leakage, entity names in the question are anonymized as “this <ENTITY_NAME>”. For instance, a final generated question asks “What brand follows this vehicle in its product line?” in Figure 3.
Language Bias Option
Language bias occurs when models overly attend to question-related information via pathways like ${Q}\rightarrow T\rightarrow A$. To assess this, we simulate conditions where only the question is provided and evaluate MLLMs’ responses. We use the answers of GPT-4 (i.e., the text-only version of GPT-4V), to ensure uniform final options across MLLMs. For example, in Figure 3, GPT-4 reponds “Tesla” to the question. Additional results using different models for generating language bias options are detailed in Section 5.3 and the prompt template is available in Appendix A.2.
Vision Bias Option
Vision bias occurs when visual information dominates (e.g., via $I\rightarrow{E}\rightarrow A$). We use the visually associated entity name (e.g., “Chrysler PT Cruiser”) as an option, to see if the model directly selects it upon encountering an option that aligns with the visual information.
Semantic Misleading Option
We introduce a semantic misleading option, such as “The Chrysler Corporation”, to challenge MLLMs’ multi-hop reasoning. This option refers to the entity that is pointed by the relation commonly owned by both the associated entity and its sampled neighbor. For example in Figure 3, upon encountering a question containing “brand” and “Chrysler PT Cruiser”, MLLMs might simply output a direct answer (e.g., “The Chrysler Corporation”), ignoring other constraints in the question (e.g., “followed by”), hence struggling to choose the correct answer (e.g., “Fiat Automobiles S.p.A.”).
Ground Truth Option
Corresponding to the causal path via ${E}\rightarrow G$ and ${S}\rightarrow G$, this option represents the final entity in the entity-relation path (e.g., “Fiat Automobiles S.p.A.”). Finally, we check and ensure that each option is unique to prevent overlap samples.
Causal Rationale Generation
Using entity-relation paths, we generate a causal rationale that aids in answering questions through a heuristic rule-based approach. As shown in Figure 3, this process begins with the associated entity and proceeds step-by-step to the ground truth. These rationales help confirm the accuracy of MLLMs’ reasoning and improve their interpretability. They also can contribute to fine-tuning MLLMs for better multi-hop reasoning.
3.3 Dataset Statistics and Quality Analysis
Statistics of Different Hops
We automatically generate training data from INFOSEEK’s train set, and development/test data from its validation set, as shown in Table 3 in Appendix. Focusing on questions with 2-hop and 3-hop depths, we set $n=1,2$ respectively. We avoid longer-hop questions to prevent potential ambiguity and complexity.
Question Distribution
We categorize the generated questions into distinct types based on their starting n-grams in Figure 15 in the Appendix. The MORE dataset showcases an extensive lexical diversity in the questions generated.
Question Quality
We analyze the lexical diversity and fluency of the generated questions, with baselines and metrics detailed in Appendix C.1. From Figure 4, MORE shows superiority in lexical diversity and fluency, even compared to human-generated datasets.
Human Evaluation
Our human evaluation confirms the high quality of generated questions and rationales, with 91% of questions and 98% of rationales deemed valid by annotators (details are in Appendix C.2).
Figure 4: Question quality of MORE compared to other VQA datasets in terms of lexical diversity and fluency.
3.4 Causal Effect Calculation on MORE
Based on the analysis in Section 2.3, we discuss how to quantify the causal effects of unimodal biases on model predictions using MORE.
Causal Effect of Questions
When intervening on questions, we keep the images constant: (1) For $\operatorname{TCE}({Q}\text{ on }A)$, we alter the question to another from the dataset pertaining to the same entity, changing both its text and core semantics, which shifts the ground truth. (2) For $\operatorname{DCE}({T}\rightarrow A)$, we let ChatGPT rephrase the question, altering its textual form but preserving its semantic meaning, resulting in no ground truth change post-intervention.
Causal Effect of Images
When intervening on images, we keep the questions constant: (1) For $\operatorname{TCE}({I}\text{ on }A)$, we replace the image with another from the dataset corresponding to the same question but featuring different entities, altering both core entity and visual context to change the ground truth post-intervention. (2) For $\operatorname{DCE}({C}\rightarrow A)$, we replace the image with another that depicts the same entity, to keep the ground truth consistent after the intervention. Besides, it is practically challenging to calculate $\operatorname{DCE}({E}\rightarrow A)$. Because intervening on core entity $E$ while maintaining the constancy of the question $Q$, visual context $C$, and ground truth answer $G$ is hard to achieve, hence not considered in further analysis.
Given the impracticality of testing all perturbations of $T$ and $C$, we randomly select 100 samples for each type of intervention and compute the average effects to determine $\operatorname{TCE}$ and $\operatorname{DCE}$ in Section 5.3. Overall, a higher $\operatorname{TCE}$ indicates better sensitivity, while a lower $\operatorname{DCE}$ indicates better robustness.
4 CAVE for Bias Mitigation
To mitigate unimodal biases in MLLMs and improve their reasoning capabilities, we propose CAVE, a causality-enhanced agent framework in this section.
For a given instance, CAVE starts by using a question decomposer to break down complex questions into simpler, step-by-step subquestions, explicitly avoiding spurious paths that may lead to incorrect answers. For each subquestion, CAVE uses a causality-enhanced reasoner to evaluate the correctness of the decomposed subquestion. Specifically, based on the aforementioned causal analysis in Section 2, it assesses whether the decomposed subquestion changes under two conditions: 1) rephrasing the question text, or 2) replacing the image with another captured by the same entity from a different angle or at a different time. If it identifies that the decomposed subquestion needs to be altered, this indicates that the previous understanding is influenced by irrelevant factors or biases, failing to capture the true semantics. In such cases, it will enforce a new round of question decomposition until it determines that the current subquestions are appropriate and accurate. Next, CAVE uses a verifier to strategically employ external tools and evaluate their outputs, acquiring the necessary context or information, such as image and text retrieval, to provide a precise answer. This iterative process of answering and verifying continues until all subquestions are resolved. By incorporating external knowledge, the final verified output integrates information from the entire reasoning process to provide a correct answer. A detailed illustration of CAVE and the prompt template is provided in Appendix E.
5 Experiments
Model LLM # Params Two-Hop, acc (%) Three-Hop, acc (%) Overall, acc (%)
Random / / 25.0 25.0 25.0
InstructBlip Vicuna 13B 17.0 16.2 16.6
mPLUG-Owl Llama 7B 12.4 11.4 11.9
LLaVA Llama 13B 20.8 13.6 17.5
Qwen-VL Qwen 7B 17.4 15.6 16.5
GPT-4v - - 17.3 16.0 16.5
GPT-4o - - 18.0 17.0 17.5
GPT-4o + CAVE - - 33.7 27.3 30.5
Gemini Pro (V) - - 25.4 20.7 22.3
Gemini Pro (V) + CAVE - - 35.6 28.8 33.2
Table 2: MLLMs’ results on the test set of MORE. We report the VQA accuracy (%) under the multi-choice settings on two-hop, three-hop, and all data, respectively.“-” denotes not released information.
5.1 Experimental Setup
Datasets
We evaluate all test data from the MORE dataset using the Multi-choice settings, where MLLMs select answers from four provided options, with a random baseline accuracy of 25%.
Baselines
We evaluate various leading MLLMs on our MORE dataset in a zero-shot fashion, including three limited-access MLLMs: GPT-4v, GPT-4o OpenAI (2023), and Gemini Pro Vision Team et al. (2023), and four open-source MLLMs: InstructBLIP (13B)  Dai et al. (2023), mPLUG-Owl (7B) Ye et al. (2023), LLaVA (v1.5, 13B) Liu et al. (2023a), and Qwen-VL (7B) Bai et al. (2023b) the details are in Appendix D. For consistent evaluation, we use standard accuracy metrics for all the models. We further quantify the causal effects of images and questions on model predictions following Section 5.3.
5.2 Evaluation Results
The results of MLLMs on MORE are shown in Table 2 and further exemplified in Appendix F. We observe that:
1) All baselines perform poorly on MORE (e.g., even the best-performing model, Gemini Pro Vision, achieves only 22.3% accuracy, which does not surpass the random baseline), indicating MLLMs’ vulnerability to biases.
2) Most models perform better on two-hop data than on three-hop data, suggesting that MLLMs’ reasoning capabilities are challenged when the problems become more complex.
3) GPT-4v falls short versus Gemini Pro Vision, possibly because we use homologous GPT-generated distractors when constructing the language bias options in Section 3.2, which particularly challenges GPT-4v’s judgment. This point is further analyzed in Section 5.3.
4) Our proposed CAVE significantly enhances GPT-4o and Gemini Pro Vision on MORE, validating the effectiveness of our method. However, the relatively low absolute values indicate ongoing challenges related to biases, suggesting the need for further research efforts.
5.3 Causal Analysis of VQA Biases
In this subsection, we select several representative MLLMs and analyze their performance through a causal lens.
Option Distribution
In Figure 5, we show the option distribution of selected MLLMs. Here we mainly utilize language bias options generated by GPT-4v and Gemini Pro Vision for comparison because they explicitly provide corresponding text-only versions. The observations are as follows:
1) Severe unimodal biases. More than 40% of the options show either language or vision bias in all models, underscoring the prevalence of unimodal biases.
2) Limited understanding ability. Models’ selection of semantically misleading options indicates some ability to combine visual and textual information, though not fully grasping the problem. This highlights the challenge our MORE dataset poses to current MLLMs.
3) Obvious selection tendency. GPT-4v often incorrectly chooses language bias options generated from GPT-4 (i.e., the text-only version of GPT-4v). Switching to Gemini Pro (i.e., the text-only version of Gemini Pro Vision) shifts this trend, with GPT-4v’s language bias selections decreasing and Gemini Pro’s increasing. These observations align with our prior analysis of GPT-4v.
Please note that discrepancies may exist between the proportions of ground truth options presented here and the accuracy values reported in Table 2, as some models’ outputs may not align with the provided option format (e.g., mPLUG-Owl), thus affecting the count of valid answers.
Figure 5: Option distribution of MLLMs.
Figure 6: Comparison of direct and total effects of image and question on prediction for MLLMs.
Causal Effects of Images and Questions
To further analyze the impact of unimodal biases on the model predictions, we assess the causal effects based on the discussion in Section 3.4. As shown in Figure 6, we select several representative MLLMs:
1) Current MLLMs exhibit high sensitivity (high $\operatorname{TCE}$), a possible reason is that instruction tuning makes models sensitive to variations in input Stolfo et al. (2023).
2) However, the model’s robustness is relatively low, as indicated by a high $\operatorname{DCE}$, meaning its predictions fluctuate even when the ground truth remains constant. This phenomenon becomes particularly pronounced when irrelevant text surface forms are introduced. This suggests that, although MLLMs can adapt well to answers, they exhibit a lack of robustness when exposed to interference, relying on spurious correlations rather than genuine causal features.
6 Related Work
Multimodal Large Language Models (MLLMs)
Recent advances in large language models (LLMs) have led to the emergence of MLLMs, which demonstrate exceptional performance in multimodal tasks OpenAI (2023); Team et al. (2023); Liu et al. (2023a); Hu et al. (2023); Yu et al. (2024). However, in contrast to the extensive evaluation of reasoning capabilities in LLMs Wei et al. (2022); Chen et al. (2024a), currently, the evaluation of MLLMs primarily emphasizes basic visual tasks Liu et al. (2023b); Fu et al. (2023); Lu et al. (2024); Chen et al. (2024b), with limited investigation into their reasoning capabilities.
Knowledge-based VQA Datasets
Existing VQA datasets Wang et al. (2017); Marino et al. (2019); Chen et al. (2023) are limited by their focus on image-related information, lack of multi-hop reasoning, open-ended answers, and reasoning rationales. They also fail to measure the effect of language and vision biases. Our MORE dataset, outlined in Table 1, addresses these shortcomings by providing a more comprehensive assessment.
Language and Vision Biases in VQA
Research reveals that some VQA models primarily depend on statistical priors from training data instead of genuinely comprehending image content Agrawal et al. (2018). These models exhibit language and vision biases; the former arises from strong correlations between specific questions and answers Abbasnejad et al. (2020); Zhu et al. (2020), and the latter from frequent co-occurrences of textual and visual elements in the dataset Si et al. (2022); Gupta et al. (2022). Recent efforts to address these biases mostly involve data augmentation Niu et al. (2021). Besides, several studies Rohrbach et al. (2018); Parcalabescu et al. (2022); Parcalabescu and Frank (2023) use examples derived from language biases to assess object hallucination in models. However, these constructed examples are typically counterfactual, with the answer types limited to yes/no, essentially judging whether the constructed statements are correct or incorrect.
7 Conclusion
This paper presents a comprehensive approach to quantifying and mitigating the unimodal biases in MLLMs. Through our causal inference framework, we provide an in-depth analysis to assess the causal effects of such biases on the model’s prediction in VQA problems. The introduced MORE dataset challenges MLLMs to engage in multi-hop reasoning and to overcome language and vision biases, thereby pushing the boundaries of their reasoning capabilities. Our proposed CAVE method demonstrates significant potential in enhancing the reasoning abilities of MLLMs.
Limitations
Our current generation of rationales is based on heuristic rules. Previous works have demonstrated the effectiveness of incorporating rationales into instructions Wei et al. (2022). Therefore, we believe that refining and polishing these rationales with an LLM (e.g., ChatGPT) could be beneficial. Besides, the Wikidata5M dataset we employed was released in 2021, and some information in the knowledge graph may be outdated. Although we have made efforts to manually verify the test set and try to ensure it does not contain incorrect information, it is still inevitable that errors may occur within the extensive training data.
Acknowledgments
We thank all the anonymous reviewers for their valuable feedback throughout the review process. This work is supported in part by Ucap Cloud and the State Key Laboratory of General Artificial Intelligence.
Appendix A Prompt Templates
A.1 Question Generation
Figure 7: Prompt template of multi-hop question generation.
We present the prompt template for generating language bias options of Section 3.2 in Figure 7.
A.2 Language Bias Option Generation
Figure 8: Prompt template of language bias option generation.
We present the prompt template for generating language bias options of Section 3.2 in Figure 8.
Appendix B Question Distribution
In Figure 15, we categorize the generated questions into distinct types, based on their starting n-grams. The dataset MORE showcases an extensive lexical diversity in the questions generated. This diversity is evidenced by variations in the introductory interrogative words (e.g., “what”, “who”, “where”, etc.), exemplified by phrases like “What is the…”, “In which country…”, and more. Such lexical richness is crucial for mitigating the vulnerability of MLLMs to linguistic variations.
Appendix C Quality Analysis Details
C.1 Question Quality
To ensure the quality of the comprising datasets, we analyze the lexical diversity and the fluency of the generated questions, which are useful for conducting a robust evaluation using questions that are linguistically diverse and coherent.
Baselines
We select extensive VQA datasets for comparison, including Visual7W Zhu et al. (2016), VQA (v2) Goyal et al. (2017), FVQA Wang et al. (2017), OK-VQA Marino et al. (2019), S3VQA Jain et al. (2021), A-OKVQA Schwenk et al. (2022), and INFOSEEK Chen et al. (2023) (contains both automated generation version and human-annotated version).
Evaluation Metrics
For lexical diversity, we utilize three metrics that are not dependent on length: moving average type-token ratio (MATTR) Covington and McFall (2010), measure of textual lexical diversity (MTLD) McCarthy (2005), and hypergeometric distribution diversity (HDD) McCarthy and Jarvis (2010). We average these three metrics for a unified assessment and employ the Lexical-Richness package Shen (2022) (version 0.5.03) for calculation. For fluency, we employ a pre-trained language model GPT2-large Radford et al. (2019) with 774M parameters to compute the perplexity of the questions, which is often used as a measure by previous work Wang et al. (2019); Cahyawijaya et al. (2021).
Figure 9: Human evaluation results of MORE.
C.2 Human Evaluation
Questions
We conduct a human evaluation of 100 questions randomly chosen from the MORE dataset to validate and assess the quality of the generated questions. This evaluation is carried out independently by three human annotators, who are provided with detailed guidelines and illustrative examples before starting the evaluation process. For each question, given the visual context and ground truth answer, we first ask two junior annotators to determine whether: 1) the question is valid, 2) the question allows for an alternative answer, 3) the question does not match the answer, or 4) the question is unclear or ambiguous. If the choices of the two annotators are inconsistent, a senior annotator checks the answers and makes the final decision. The average inter-annotator agreement is 88.6% (Cohen’s kappa).
Rationales
We also conduct a human evaluation of the causal rationales, following the same procedure as described above. The difference is that here we provide annotators with only two options: to assess whether the generated rationales are valid.
As shown in Figure 9, the results are encouraging, with 91% questions and 98% rationales being classified as valid by the annotators, further demonstrating the quality of our datasets.
Appendix D Baselines
For open-source MLLMs, we consider the following baselines:
1) InstructBLIP Dai et al. (2023), an instruction-tuned version of BLIP-2 on various tasks including VQA. We employ its InstructBLIP-Vicuna Chiang et al. (2023)-13B variant.
2) mPLUG-Owl Ye et al. (2023), which proposes a new two-stage training method for aligning images and text. We employ its mPLUG-Owl-Llama Touvron et al. (2023a)-7B variant.
3) LLaVA Liu et al. (2023a), which translates images into texts of captions and bounding boxes, and prompts GPT-4 to generate a multimodal instruct-tuning dataset in the context of seed examples. We employ its LLaVA-Llama Touvron et al. (2023a)-13B variant.
4) Qwen-VL Bai et al. (2023b), which builds upon Qwen Bai et al. (2023a) and employ multi-stage training pipeline. Qwen-VL facilitates grounding and text comprehension by aligning image-caption-box tuples. It processes inputs of images, text, and bounding boxes, and produces corresponding text and bounding boxes as outputs.
Appendix E Details of the CAVE Framework
In this section, we first present the overall framework as shown in Figure 10, and then we will go over each part of it in detail.
E.1 Overall Framework
Given a question $Q$ and an image $I$, the VQA task demands the system to return an output $A$ that concisely answers the question. As shown in Figure 10, we first initialize a question decomposer $D$ to analyze $Q$ and break it down into manageable subquestions. For each subquestion, we introduce a causality-enhanced reasoner $C$ to evaluate the correctness of the decomposed subquestion. Then, we employ a verifier $V$ to confirm the accuracy of the original answer to each subquestion. The generation verifier typically involves active information-seeking and answer-verification, which acquires the necessary context or information needed to investigate and revise the answers. This includes two optional operations: image retrieval to seek images similar to $I$ and determine their titles, or text retrieval with a specific query to fetch pertinent documents and summarize their content. This iterative process of answering and verifying will continue until we resolve each subquestion. Finally, the verified answer $A$ is output following the aforementioned reasoning process and retrieved information.
Figure 10: An overview of our proposed CAVE Framework.
E.2 Question Decomposer
For a given VQA problem, MLLMs may simply exploit a spurious shortcut to make predictions. In order to alleviate this issue, motivated by Chain-of-Thought reasoning Wei et al. (2022), we encourage MLLMs to decompose the question $Q$ before outputting the answer, so as to gradually solve a complex question that requires multi-hop reasoning. As shown in Figure 10, for the question “Which country hosted the next World Cup after this venue?”, our decomposer breaks it down into two subquestions :
  • 1.
    What this venue is?
  • 2.
    Which country hosted the next World Cup after Allianz Arena?
Such decomposition will explicitly constrain the model from comprehending and extracting the truth semantics of the question, thus avoiding simply exploring a spurious path to give the answer.
E.3 Causality-enhanced Reasoner
The accuracy of sub-question decomposition plays a crucial role in the subsequent reasoning process. To address this, we introduce causality-enhanced self-reflection, which explicitly guides the model in validating the sub-question it has decomposed. Specifically, based on the proposed causal framework, we guide the model to assess whether the decomposed subquestion changes under two conditions: 1) rephrasing the question text, or 2) replacing the image with another captured by the same entity from a different angle or at a different time. If the model identifies that the decomposed subquestion needs to be altered, this indicates that its initial understanding is influenced by irrelevant factors or biases, failing to capture the true semantics. In such cases, we will enforce a re-decomposition of the question.
E.4 Verifier
Some works have found that vision illusion and language hallucination may appear in the process of MLLMs’ response generation Guan et al. (2023); Yang et al. (2023). To alleviate this issue, we adopt the retrieval-augmented generation approaches Khandelwal et al. (2020); Chen et al. (2022). Specifically, we consider two different retrieval ways for the verifier to choose during each verification step: image retrieval and text retrieval.
Image Retrieval
Although our framework is applicable to any image retrieval method, in this paper, we mainly utilize Google Image Search to obtain a broad range of information related to the image as provided by Google Lens API [^1] Web interface available at https://images.google.com.. This information encompasses various details, such as knowledge graph entities and captions of analogous or identical images. The availability of these details can vary based on the image input provided to Google Image Search. Then, the verifier gleans relevant information from captions associated with visually similar images, so as to verify the original answer and conduct the next round of reasoning.
Text Retrieval
Similarly, our framework is applicable to any text retrieval method, we explore a simple, off-the-shelf dense retriever for Wikipedia, GTR Karpukhin et al. (2020), as our text retriever. First, the verifier constructs a query to perform text retrieval according to the currently generated context, and then the query is input into a GTR model to get related document titles and contents. Finally, the verifier will fetch pertinent documents and summarize their content to verify the immediate answer.
E.5 Final Verified Answer
Finally, the improved response that takes verification into account is generated. This is executed by a final prompt where the context takes into account all of the previous reasoning steps, the baseline response, and the verification question-answer pairs, so that the corrections can take place.
Appendix F Case Study
We conduct a case study on the development set of MORE in Figure 11$\sim$14, including both the “Open-ended” and “Multi-Choice” settings.
Figure 11: Two-hop case in the dev set of MORE. The wrong answers are marked in red and the correct is in green.
Figure 12: Two-hop case (multi-choice setting) in the dev set of MORE. The wrong answers are marked in red.
Figure 13: Three-hop case in the dev set of MORE. The wrong answers are marked in red.
Figure 14: Three-hop case (multi-choice setting) in the dev set of MORE. The wrong answers are marked in red and the correct is in green.
Figure 15: Question prefix distribution in MORE. The arc length is proportional to the number of questions containing the word.